The balances of two investments are changing at steady rates. One investment of $1450 is decreasing by $105.75 each month. The other investment of $825 is increasing by $144.25 each month. Write and solve an equation to find how many months it takes the balances of the investments to be the same. Let m represent the number of months.
Answer by mananth(16946)

The balances of two investments are changing at steady rates. One investment of $1450 is decreasing by $105.75 each month. The other investment of $825 is increasing by $144.25 each month. Write and solve an equation to find how many months it takes the balances of the investments to be the same. Let m represent the number of months.
One investment of $1450 is decreasing by $105.75 each month.
B1= 1450-(105.75m)
The other investment of $825 is increasing by $144.25 each month.
B2=825+144.25m
when will B1=B2
1450-105.75m=825+144.25m
Rearrange
1450-825=144.25m+105.75m
625=250m
m=625/250=2.5
Balances of the two investments will be the same at 2.5 month